# Hungarian Research Network (HUN-REN) Laboratory on Engineering and Management Intelligence Research Group of Operations Research and Decision Systems

### v11

@@ Abstract
Several sports tournaments contain a round-robin group stage where the teams are assigned to groups subject to some constraints. Hence, the organisers usually use a computer-assisted random draw to avoid any dead end, a situation when the teams still to be drawn cannot be assigned to the remaining empty slots. However, this procedure is known to be unfair: the feasible allocations are not equally likely. We quantify the unfairness of the 2018 FIFA World Cup draw and evaluate its effect on the probability of qualification for the knockout stage for each national team. The official draw order of Pot 1, Pot 2, Pot 3, Pot 4 turns out to be a significantly better option than the 23 other draw orders with respect to the unwanted distortions. Nonetheless, the non-uniform draw distorts the probability of qualification by more than one percentage point for two countries. Our results call attention to the non-negligible role of draw order and make it possible for policymakers to decide whether the use of fairer draw mechanisms is justified.
@@ 1 Introduction
Until 2014, the FIFA World Cup draw divided the teams into pots mainly according to their geographic area, which caused serious unfairness in certain years such as in 1990 (Jones,, 1990), 2006 (Rathgeber and Rathgeber,, 2007), and 2014 (Guyon,, 2015). In particular, the top teams had different chances of being placed in a group with difficult opponents. Consequently, Guyon, (2014) has recommended three options to create balanced and geographically diverse groups—and one of them has been adopted for both the 2018 and 2022 FIFA World Cups (Guyon,, 2018; Csató, 2023a, ).
However, this procedure, based on a computer-assisted random draw to avoid any dead end (a situation when the teams still to be drawn cannot be assigned to the remaining empty slots), is not uniformly distributed: the valid assignments are not equally likely to occur (Roberts and Rosenthal,, 2023). Hence, Roberts and Rosenthal, (2023) propose two unbiased mechanisms that use balls and bowls, making them suitable for a televised draw. On the other hand, they require random simulations, which might threaten transparency.
@@
The current work aims to fill this research gap. In particular, the 2018 FIFA World Cup has been chosen as a case study since the 2022 FIFA World Cup draw has contained an inherent bias due to the uncertainty in the set of teams qualified (Csató, 2023a, ), and the teams qualifying for the 2026 FIFA World Cup are naturally unknown at the time of writing.
@@
Our topic is hugely relevant to sport policy. In an ideal environment, the rules governing a sport prevent any player or team from gaining an unfair advantage (Arnold,, 1994). But, if the groups are created to promote attractiveness and uniqueness by maximising the number of intercontinental games (Guyon,, 2015), the sequential draw method of FIFA may threaten fairness, while the alternative solutions remain more complex and less transparent (Roberts and Rosenthal,, 2023). Therefore, policymakers could maintain the balance between fairness and integrity only by investigating different fairness indicators that will be calculated in the following (Scelles et al.,, 2024).
@@ 2 Related literature
Guyon, (2014) presents alternative proposals to retain the practicalities of the FIFA World Cup draw but improve its outcome. One of them has been adopted by FIFA for the 2018 World Cup draw (Guyon,, 2018). Another recommended solution can be compared to the flawed FIFA rule at https://www.nytimes.com/interactive/2014/06/03/upshot/world-cup-draw-simulation.html.
@@
Other studies deal with the UEFA Champions League where a constrained draw mechanism has been used in the Round of 16 between the 2003/04 and 2023/24 seasons. However, this method is distinct from the FIFA World Cup draw procedure (Csató, 2023b, ). Wallace and Haigh, (2013) verify that the possible assignments are not equally likely and highlight the connection of the UEFA draw mechanism to Hall’s marriage theorem, see also Haigh, (2019, Section 3.6). Kiesl, (2013) computes the bias in the 2012/13 season and outlines some fair—but uninteresting to watch—methods. According to Klößner and Becker, (2013), the draw system inherently implies different probabilities for certain assignments, which are translated into more than ten thousand Euros in expected revenue due to the substantial amount of prize money. Finally, Boczoń and Wilson, (2023) reveal how the UEFA draw procedure affects expected assignments and address the normative question of whether a fairer randomisation mechanism exists. The current design is verified to come quantitatively close to a constrained best in fairness terms.
@@ 4.1 The effect of draw order
However, an appropriate relabelling of the pots may bring the FIFA World Cup draw closer to the uniform distribution. Therefore, all possible orders of the pots are examined such that the pre-assignment of Russia to Group A is retained. Since the teams can be drawn in $7!\times\left(8!\right)^{3}\approx 3.3\times 10^{17}$ different orders, it is almost impossible to derive exact theoretical results. Consequently, the 25 draw mechanisms—the draw procedure of FIFA with the 24 possible orders of the pots and the uniformly distributed rejection mechanism (Roberts and Rosenthal,, 2023)—are analysed on the basis of 1 million randomly generated draw orders, analogous to (Roberts and Rosenthal,, 2023).
@@
$\Delta_{ij}$ $=$ $P\left(\text{teams $i$ and $j$ are assigned to the same group bymechanism $M$}\right)-$
[…]
$p_{ij}^{M}-p_{ij}^{U}.$
However, when the biases are aggregated for the teams, we will always sum their absolute value $\lvert\Delta_{ij}\rvert$.
@@ 4.2 The consequences of non-uniform distribution
The same simulation methodology has recently been applied in some studies on tournament design (Csató, 2022b, ; Csató, 2023a, ; Stronka,, 2024). Note that even if this Elo-based approach is not necessarily the best available simulation model, it is only used to compare the outcomes under the uniform and non-uniform draw. Thus, our assumptions seem to be adequate for comparative purposes as emphasised by Appleton, (1995).
@@ 5 Discussion
  • The runners-up were drawn first in the UEFA Champions League Round of 16 (Boczoń and Wilson,, 2023; Klößner and Becker,, 2013), ans the unseeded teams were drawn first in the UEFA Europa League Round of 32 (Csató, 2022a, );
@@ Acknowledgements
This paper could not have been written without my father (also called László Csató), who has coded the simulations in Python.
Eleven anonymous reviewers provided valuable remarks and suggestions on earlier drafts.

### v12

@@ Abstract
Several sports tournaments contain a round-robin group stage where the teams are assigned to groups subject to some constraints. Hence, the organisers usually use a computer-assisted random draw to avoid any dead end, a situation when the teams still to be drawn cannot be assigned to the remaining empty slots. This procedure is known to be unfair: the feasible allocations are not equally likely. We quantify the unfairness of the 2018 FIFA World Cup draw and evaluate its effect on the probability of qualification for the knockout stage for each national team. The official draw order of Pot 1, Pot 2, Pot 3, Pot 4 turns out to be a significantly better option than the 23 other draw orders with respect to the unwanted distortions. Nonetheless, the non-uniform draw distorts the probability of qualification by more than one percentage point for two countries. Our results call attention to the non-negligible role of draw order and make it possible for policymakers to decide whether the use of fairer draw mechanisms is justified.
@@ 1 Introduction
Until 2014, the FIFA World Cup draw divided the teams into pots mainly according to their geographic area, which caused serious unfairness in certain years such as in 1990 (Jones,, 1990), 2006 (Rathgeber and Rathgeber,, 2007), and 2014 (Guyon,, 2015). In particular, the top teams had different chances of being placed in a group with difficult opponents. Consequently, Guyon, (2014) has recommended three options to create balanced and geographically diverse groups, which has inspired FIFA to change the draw procedure used in the 2018 and 2022 FIFA World Cups (Guyon,, 2018; Csató,, 2023).
However, this mechanism, based on a computer-assisted random draw to avoid any dead end (a situation when the teams still to be drawn cannot be assigned to the remaining empty slots), is not uniformly distributed: the valid assignments are not equally likely to occur (Roberts and Rosenthal,, 2023). Hence, Roberts and Rosenthal, (2023) propose two unbiased mechanisms that use balls and bowls, making them suitable for a televised draw. On the other hand, they require random simulations, which might threaten transparency.
@@
The current work aims to fill this research gap. In particular, the 2018 FIFA World Cup has been chosen as a case study since the 2022 FIFA World Cup draw has contained an inherent bias due to the uncertainty in the set of teams qualified (Csató,, 2023), and the teams qualifying for the 2026 FIFA World Cup are naturally unknown at the time of writing.
@@
Our topic is hugely relevant to governing bodies in sports. In an ideal environment, the rules governing a sport prevent any player or team from gaining an unfair advantage (Csató,, 2021; Devriesere et al.,, 2024). But the sequential draw method of FIFA may threaten fairness, while the alternative solutions remain more complex and less transparent (Roberts and Rosenthal,, 2023). Therefore, policymakers could maintain the balance between fairness and integrity only by investigating different fairness indicators (Scelles et al.,, 2024).
@@ 2 Related literature
Guyon, (2014) presents alternative proposals to retain the practicalities of the FIFA World Cup draw but improve its outcome. One of them can be compared to the flawed FIFA rule at https://www.nytimes.com/interactive/2014/06/03/upshot/world-cup-draw-simulation.html.
@@
Other studies deal with the UEFA Champions League where a constrained draw mechanism has been used in the Round of 16 between the 2003/04 and 2023/24 seasons. However, this method is distinct from the FIFA World Cup draw procedure (Csató,, 2024). Wallace and Haigh, (2013) verify that the possible assignments are not equally likely and highlight the connection of the UEFA draw mechanism to Hall’s marriage theorem, see also Haigh, (2019, Section 3.6). Kiesl, (2013) computes the bias in the 2012/13 season and outlines some fair—but uninteresting to watch—methods. According to Klößner and Becker, (2013), the draw system inherently implies different probabilities for certain assignments, which are translated into more than ten thousand Euros in expected revenue due to the substantial amount of prize money. Finally, Boczoń and Wilson, (2023) reveal how the UEFA draw procedure affects expected assignments and address the normative question of whether a fairer randomisation mechanism exists. The current design is verified to come quantitatively close to a constrained best in fairness terms.
@@ 4.1 The effect of draw order
However, an appropriate relabelling of the pots may bring the FIFA World Cup draw closer to the uniform distribution. Therefore, all possible orders of the pots are examined such that the pre-assignment of Russia to Group A is retained. Since the teams can be drawn in $7!\times\left(8!\right)^{3}\approx 3.3\times 10^{17}$ different orders, it is almost impossible to derive exact theoretical results. Consequently, the 25 draw mechanisms—the draw procedure of FIFA with the 24 possible orders of the pots and the uniformly distributed rejection mechanism (Roberts and Rosenthal,, 2023)—are analysed on the basis of 1 million randomly generated draw orders, analogous to Roberts and Rosenthal, (2023).
@@
$\Delta_{ij}=p_{ij}^{M}-p_{ij}^{U},$
where $p_{ij}^{M}$ and $p_{ij}^{U}$ are the probabilities that teams $i$ and $j$ are assigned to the same group by a mechanism $M$ and the uniform draw, respectively. When the biases are aggregated for the teams, we will always sum their absolute value $\lvert\Delta_{ij}\rvert$.
@@ 4.2 The consequences of non-uniform distribution
The same simulation methodology has recently been applied in some studies on tournament design (Csató, 2022b, ; Csató,, 2023; Stronka,, 2024). Note that even if this Elo-based approach is not necessarily the best available simulation model, it is only used to compare the outcomes under the uniform and non-uniform draw. Thus, our assumptions seem to be adequate for comparative purposes as emphasised by Appleton, (1995).
@@ 5 Discussion
  • The runners-up were drawn first in the UEFA Champions League Round of 16 (Boczoń and Wilson,, 2023; Klößner and Becker,, 2013), and the unseeded teams were drawn first in the UEFA Europa League Round of 32 (Csató, 2022a, );
@@ Acknowledgements
This paper could not have been written without my father (also called László Csató), who has primarily coded the simulations in Python.
Eleven colleagues and anonymous reviewers provided valuable remarks and suggestions on earlier drafts.